# ReAct Meets ActRe: Autonomous Annotation of Agent Trajectories for Contrastive Self-Training

### v2

@@ 3.1 AlfWorld
Tables 1 and 2 show the performance comparison of our framework. For the single trial setting, the overall success rate of our agent reaches $96\%$ at $2$-nd round and matches the prior SoTA (LM-BUTLER). However, our agent is trained with a QLoRA of 26M parameters and 600 training tasks, while LM-BUTLER is fine-tuned from a full GPT2-medium of 355M parameters and all 3,553 training tasks. For the iterative refinement setting, our agent obtains 100% success by accumulating the decision-making trials of all the $4$ trained agents from each round. The accumulated trajectory set accounts for the significance of the performance. Table 3 shows that the success rate of the trajectories composed by the agent on the training tasks improves continually. More experimental results are covered in Appendix C.1.
@@ C.1 AlfWorld
Test (in-distribution) 35 27 16 25 13 24 140
Test (out-of-distribution) 24 31 23 21 18 17 134
Table 13: The statistics of tasks used in training, validation, and testing. : For all experimental results except for Table 14, we report the success rate on the 134 out-of-distribution test scenarios following the settings in previous work.
Tables 14 and 15 exhibit more experimental results on the in-distribution testing tasks and the validation tasks.
@@
Ours (Round=0) 86 67 94 80 85 75 80
Ours (Round=1) 91 89 88 96 92 67 87
Ours (Round=2) 97 89 100 88 92 79 91
Ours (Round=3) 100 96 100 96 77 79 93
@@
Ours (Round=0) 11/12   9/10 5/7 11/14 7/8 9/9 52/60
Ours (Round=1) 11/12 10/10 6/7 13/14 7/8 8/9 55/60
Ours (Round=2) 12/12 10/10 7/7 13/14 8/8 8/9 58/60
Ours (Round=3) 12/12 10/10 7/7 12/14 8/8 7/9 56/60

### v3

# ReAct Meets ActRe: When Language Agents Enjoy Training Data Autonomy

@@ 3.1 AlfWorld
Tables 1 and 2 show the performance comparison of our framework. For the single trial setting, the overall success rate of our agent reaches $96\%$ at $2$-nd round and matches the prior SoTA (LM-BUTLER). However, our agent is trained with a QLoRA of 26M parameters and 600 training tasks, while LM-BUTLER is fine-tuned from a full GPT2-medium of 355M parameters and all 3,553 training tasks. Besides, our agent demonstrates constant performance improvements with the $4$ rounds in the held-out seen evaluation scenarios and outperforms LM-BUTLER (Table 14 in Appendix C.1). For the iterative refinement setting, our agent obtains 100% success by accumulating the decision-making trials of all the $4$ trained agents from each round. The accumulated trajectory set accounts for the significance of the performance. Table 3 shows that the success rate of the trajectories composed by the agent on the training tasks improves continually. More details are covered in Appendix C.1.
@@ C.1 AlfWorld
Test (held out, seen) 35 27 16 25 13 24 140
Test (held out, unseen) 24 31 23 21 18 17 134
Table 13: The statistics of tasks used in training, validation, and testing. : For all experimental results except for Table 14, we report the success rate on the 134 held-out unseen test scenarios following the settings in previous work.
Tables 14 and 15 exhibit more experimental results on the 140 held-out seen testing tasks (official split) and the 60 validation tasks (our use).
@@
A3T (Round=0) 86 67 94 80 85 75 80
A3T (Round=1) 91 89 88 96 92 67 87
A3T (Round=2) 97 89 100 88 92 79 91
A3T (Round=3) 100 96 100 96 77 79 93
@@
A3T (Round=0) 11/12   9/10 5/7 11/14 7/8 9/9 52/60
A3T (Round=1) 11/12 10/10 6/7 13/14 7/8 8/9 55/60
A3T (Round=2) 12/12 10/10 7/7 13/14 8/8 8/9 58/60
A3T (Round=3) 12/12 10/10 7/7 12/14 8/8 7/9 56/60